OpenStudy (anonymous):

what is the perimeter of a rectangle that has a length of 6 inches and a width of 3 inches

OpenStudy (anonymous):

a rectangle as two equal opposite sides. so the Perimeter = 2L + 2W P = 2 (6) + 2(3) = 18 +6 = 24